# Wenfan Ou and Sheng Bi

### v3

@@ 7.2 Online and Bayesian Regime
Another significant drawback of traditional RMDPs lies in their inability to “learn” from uncertainty, which easily causes overly conservative policies due to the extensive uncertainty sets. To tackle this challenge and infuse the adaptive learning characteristic of RL, there is a growing trend to incorporate online learning or Bayesian concepts into the design of uncertainty/ambiguity sets, Bellman updates, environmental configurations, and other facets of RMDPs. Such efforts are expanding the scope of RMDPs and injecting fresh dynamism into the field.
@@
Out of consideration for limiting the size of uncertainty set, Lim et al. (2013) assume there is a unique but unknown subset $\mathcal{Z}$ of state–action pairs behaving adversarially while all others are stochastic, as in non-robust MDPs. In essence, they propose a two-step framework: divide pairs into two categories, then compute respectively. Motivated by Bubeck & Slivkins (2012) that investigate a similar scenario in a multi-armed bandit setting, the authors implement a statistical hypothesis test, termed consistency check, after executing $(s,a)$ every time. All state-action pairs that pass the consistency check are stochastic, while the pairs that do not pass the check will consist of $\mathcal{Z}$. Combining the consistency check with UCRL2, a well-known online algorithm that treats every state-action pair as stochastic (Auer et al. 2009), the authors develop two algorithms for finite-horizon and infinite-horizon cases, respectively. Adhering to the optimistic property of UCRL2, the results of the proposed algorithms are less conservative than the standard RMDPs.
@@
Sharing similar concerns but adopting distinct methodologies, Petrik & Russel (2019) observe that reliance on concentration inequalities results in conservative ambiguity sets, and then the authors introduce Bayesian inference to construct tight ambiguity sets. In the paper, the authors focus on a safe return measured by Value-at-Risk (VaR) and aim to improve the lower bound when the value function is unknown. Under Bayesian assumptions, the transition probabilities $\mathbf{p}_{sa}$ is a random variable with a prior distribution. Specifically, they propose the robustification with sensible value functions (RSVF) concept, in which the ambiguity set $\mathcal{P}_{sa}$ is derived from a set of possible value functions $\mathcal{V}$. Meanwhile, given any $v$, they approximate the optimal ambiguity set as
$\mathcal{K}_{sa}(v)=\left\{\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|}\left|\mathbf{p}_{sa}^{\top}v\leq\operatorname{VaR}_{\mathbf{p}^{0}}^{\theta}[\mathbf{p}_{sa}^{0\top}v]\right\}\right.,$ (35)
where $\mathbf{p}^{0}$ serves as the nominal probability.
@@
As augmenting value function into $\mathcal{V}$, the algorithm updates $\mathcal{P}_{sa}$ and currently optimal value function $\hat{v}$. If the intersection $\mathcal{K}_{sa}(\hat{v})\cap\mathcal{P}_{sa}\neq\emptyset$ for all $(s,a)$ pairs, the algorithm deems the current policy and value to be safe. For theoretical contributions, they provide a finite iteration probabilistic guarantee that is beneficial from their special formulation and empirically compare the safe estimates from RSVF and other non-Bayesian methods. Empirical results demonstrate that RSVF indeed provides much tighter ambiguity sets and safe returns.
@@
Contrary to most studies that concentrate on developing innovative robust algorithms to learn from the original environment, Wang, Gadot, Kumar, Levy & Mannor (2023) introduce an RL framework designed to approximate an adversarial environment and generate adversarial data for training purposes. This approach is inspired by Kumar et al. (2023) which reveals that the adversarial distribution fundamentally alters the next-state transition probability from the nominal one. Compared to applying intricate iterations of robust methods, directly employing non-robust methods in adversarial environments will be more computationally efficient, thus addressing scalability challenges in RMDPs with high-dimensional domains. Different from the $L_{p}$-norm ambiguity sets used by Kumar et al. (2023), the authors construct KL-divergence ambiguity sets under $(s,a)$-rectangular assumption, which circumvents the issue of different supports between nominal and adversarial distribution. They also derive a closed-form expression for the adversarial distribution, depicted as a re-weighted nominal distribution. Consequently, the framework can be straightforwardly implemented by simulating the next state under the nominal dynamic and choosing it with corresponding adversarial weight. After realizing the next state, $(s,a,s^{\prime})$ tuple is added to the data buffer, and the policy is trained with data from the buffer via any non-robust RL method.
@@ 7.3 Risk, Regularization, and Robustness
Our pursuit of robustness aims to enhance the generalization performance and mitigate the uncertainty when the underlying truth is ambiguous. This concept resonates across various fields, e.g., finance and machine learning (ML), where both risk-aware and regularized formulations have seen considerable success. Notably, a body of literature attempts to bridge these concepts with RO and DRO (Shapiro 2017, Namkoong & Duchi 2017, Shapiro 2021, Shafieezadeh-Abadeh et al. 2019, Blanchet & Murthy 2019), and also RMDPs. These studies offer diverse interpretations of robustness, revealing connections that shed light on the models of Markov Decision Processes (MDPs). This understanding helps to partially alleviate the conservatism inherent in robust policies.
In terms of risk-aware MDPs, Osogami (2012) explores the equivalence between RMDPs and risk-aware MDPs with an expected exponential utility objective. By leveraging the properties of this objective, the author reformulates risk-aware MDPs into a general RMDP that aims to minimize the expected cumulative cost adjusted by the KL-divergence penalty. This penalty quantifies the discrepancy between the nominal and the worst-case distribution. Building on this foundation, Bäuerle & Glauner (2022) broaden these equivalence findings to encompass all risk-aware MDPs that employ spectral risk measures, which are known as a class of coherent risk measures like the Expected Shortfall. However, the risk measures examined are confined to the Markov risk measure (Ruszczyński 2010), known to be potentially non-gradient-dominant (Huang et al. 2021).
@@
where $Q_{\tau}(X):=\inf\{x\mid P(X\leq x)\geq\tau\}$ and $\tau\in(0,1)$ is the quantile value. The authors reformulate (38) as a max-min formulation and have pioneered the verification that the quantile-based value-to-go function is tantamount to solving a specific optimization problem for a given action. This breakthrough facilitates the application of dynamic programming techniques. Additionally, they craft efficient algorithms and conduct a detailed complexity analysis. In terms of model-free approaches, Yu & Shen (2022) adapt CVaR as a special one-step conditional risk measure to preserve time consistency and subsequently reformulate the risk-averse MDP as a risk-neutral counterpart with augmented action space and the adjusted immediate rewards. Furthermore, the authors show that the Bellman operator in this formulation is a contraction mapping, and consequently, the authors extend Q-learning results into this case and develop a risk-averse deep Q-learning algorithm.
The concept of weaving regularization into dynamic programming also has been thoroughly investigated within the reinforcement learning field. (Kaufman & Schaefer 2013, Neu et al. 2017, Kostrikov et al. 2021, Ho et al. 2021, Kumar et al. 2022). While these regularized variants of value iteration and policy iteration have been developed, Grand-Clément & Petrik (2022) first propose the convex optimization formulation of RMDPs. This innovative approach integrates an entropic regularization of the robust Bellman operator with a change of variables involving exponential and logarithmic functions. By conducting this convex reformulation, arbitrary algorithms designed for solving convex optimization problems can be incorporated with the reformulated RMDPs. This fresh perspective has paved the way for the development of innovative algorithms tailored to RMDPs.
@@ 8 Conclusion
This survey provides an overview of recent methodologies for modeling RMDPs with transition uncertainty models. We start with a framework that encompasses three primary means of utilizing information to construct ambiguity sets in RMDPs. The parametric approach is the simplest yet most restrictive way to characterize the uncertainty, usually requiring constrained support or distribution family. Moment-based ambiguity sets are more suitable to exogenous randomness and concrete problem background. As the most popular approach, discrepancy-based RMDPs match in a data-driven manner. All three mainstream discrepancy measures–norm, $\phi$-divergence, and Wasserstein distance, can determine the size of ambiguity sets that contain underlying probability measures or distributions with high probability. It is noteworthy that norm and $\phi$-divergence may be more favored in fast algorithm research, while Wasserstein distance displays more powerful performance in modeling and practice.
We also review recent efforts that depart from rectangular assumptions. Rectangularity is the key assumption in most RMDPs research, which allows decomposability to ensure that RMDPs can be solved in polynomial time. We summarize three class definitions of rectangularity and prove that solving non-rectangular RMDPs is NP-hard. There are some attempts to relax traditional rectangular assumptions. While some attempts have been made to relax the traditional rectangular assumption, these coupled uncertainty modeling approaches often compromise the flexibility of the proposed model. Besides the minimax framework, we explore related paradigms in the literature that seek to achieve similar performance to traditional RMDPs. These new paradigms in RMDPs are often represented as combinations of RL and other techniques. These works establish a bridge between RMDPs and other areas, bringing new perspectives for hedging uncertainty.

### v4

@@ 7.2 Online and Bayesian Regime
As the focus of MDPs or RMDPs is planning, another significant drawback of traditional RMDPs lies in their inability to “learn” from uncertainty, which easily causes overly conservative policies due to the extensive uncertainty sets. To tackle this challenge and infuse the adaptive learning characteristic of RL, there is a growing trend to incorporate online learning or Bayesian concepts into the design of uncertainty/ambiguity sets, Bellman updates, environmental configurations, and other facets of RMDPs. Such efforts are expanding the scope of RMDPs and injecting fresh dynamism into the field.
@@
Out of consideration for limiting the size of uncertainty set, Lim et al. (2013) assume there is a unique but unknown subset $\mathcal{Z}$ of state-action pairs behaving adversarially while all others are stochastic, as in non-robust MDPs. In essence, they propose a two-step framework: divide state-action pairs into two categories first and then handle them respectively. Motivated by Bubeck & Slivkins (2012) that investigate a similar scenario in a multi-armed bandit setting, the authors employ a statistical hypothesis test, termed consistency check, after executing arbitrary state-action pair $(s,a)$ every time. All state-action pairs that pass the consistency check are stochastic, which means the incurred rewards and transitions are safely accepted, while the pairs that do not pass the check will be included in $\mathcal{Z}$. Combining the consistency check with UCRL2, a well-known online algorithm that treats every state-action pair as stochastic (Auer et al. 2009), the authors develop two algorithms for finite-horizon and infinite-horizon cases, respectively. Following the optimistic property of UCRL2, the results of the proposed algorithms are less conservative than the standard RMDPs.
@@
Sharing similar concerns but adopting distinct methodologies, Petrik & Russel (2019) observe that reliance on concentration inequalities results in conservative uncertainty sets. To circumvent this issue, the authors propose leveraging Bayesian inference to construct tighter uncertainty sets. They focus on a safe return measured by Value-at-Risk (VaR) and aim to improve the lower bound when the value function is unknown. Under Bayesian assumptions, the transition probabilities $\mathbf{p}_{sa}$ is a random variable with a prior distribution. Specifically, they introduce the robustification with sensible value functions (RSVF) concept, in which the uncertainty set $\mathcal{P}_{sa}$ is derived from a set of possible value functions $\mathcal{V}$. For any given value function $v\in\mathcal{V}$, they approximate the optimal uncertainty set $\mathcal{K}_{sa}(v)$ as
$\mathcal{K}_{sa}(v)=\left\{\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|}~{}\left|~{}\mathbf{p}_{sa}^{\top}v\leq\operatorname{VaR}_{\hat{\mathbf{p}}}^{\theta}[\hat{\mathbf{p}}_{sa}^{0\top}v]\right\}\right.,$ (57)
where $\hat{\mathbf{p}}$ serves as the nominal probability. This set essentially defines a set of plausible transition probabilities $\mathbf{p}_{sa}$ whose expected value meets the VaR criterion for being “sufficiently safe” regarding the current value function $v$.
@@
Note that $\mathcal{K}_{sa}(v)$ is sufficient to provide a reasonable uncertainty set if the value function $v$ is known. However, in the context of this paper, the value function is unknown. As new value functions are progressively added to $\mathcal{V}$, RSVF updates the uncertainty set $\mathcal{P}_{sa}$ to append plausible $\mathbf{p}_{sa}$ that is “close” enough to the worst-case distribution within the current $\mathcal{K}_{sa}(v),v\in\mathcal{V}$. When $\mathcal{P}_{sa}$ is updated, we can derive the (currently) optimal value function $\hat{v}^{*}$ and policy $\hat{\pi}^{*}$. If the intersection $\mathcal{K}_{sa}(\hat{v}^{*})\cap\mathcal{P}_{sa}\neq\emptyset$ for all $(s,a)$ pairs, which means that $\hat{v}^{*}$ is “sufficiently robust” for some adversarial distributions, the algorithm deems the current policy $\hat{\pi}^{*}$ to be safe. They provide a finite iteration probabilistic guarantee and empirically compare the safe estimates from RSVF and other non-Bayesian methods. Empirical results demonstrate that RSVF indeed provides much tighter ambiguity sets and safe returns.
@@
Contrary to most studies that concentrate on developing innovative robust algorithms to learn from the original environment, Wang et al. (2023) introduce an RL framework designed to approximate an adversarial environment and generate adversarial data for training purposes. This approach is inspired by Kumar et al. (2023) which reveals that the adversarial distribution fundamentally alters the next-state transition probability from the nominal one. Compared to applying intricate iterations of robust methods, directly employing non-robust methods in adversarial environments will be more computationally efficient, thus addressing scalability challenges in RMDPs with high-dimensional domains. Different from the $L_{p}$-norm ambiguity sets used by Kumar et al. (2023), the authors construct KL-divergence ambiguity sets under $(s,a)$-rectangular assumption, which circumvents the issue of different supports between nominal and adversarial distribution. They also derive a closed-form expression for the adversarial distribution, depicted as a re-weighted nominal distribution. Consequently, the framework can be straightforwardly implemented by simulating the next state under the nominal dynamic and choosing it with corresponding adversarial weight. After realizing the next state, $(s,a,s^{\prime})$ tuple is added to the data buffer, and the policy is trained with data from the buffer via any non-robust RL method.
@@ 7.3 Risk, Regularization, and Robustness
Our pursuit of robustness aims to enhance performance and mitigate uncertainty when the underlying truth is ambiguous. This concept resonates across various fields, e.g., finance and machine learning (ML), where both risk-aware and regularized formulations have seen considerable success. Notably, a body of literature attempts to bridge these concepts with RO and DRO (Shapiro 2017, Namkoong & Duchi 2017, Shapiro 2021, Shafieezadeh-Abadeh et al. 2019, Blanchet & Murthy 2019), and also RMDPs. These studies offer diverse interpretations of robustness, revealing connections that shed light on the models of MDPs. This understanding helps to partially alleviate the conservatism inherent in robust policies.
In terms of risk-aware MDPs, Osogami (2012) explores the equivalence between RMDPs and risk-aware MDPs with an expected exponential utility objective. By leveraging the properties of this objective, the author reformulates risk-aware MDPs into a general RMDP that aims to minimize the expected cumulative cost adjusted by the KL-divergence penalty. This penalty quantifies the discrepancy between the nominal and the worst-case distribution. Building on this foundation, Bäuerle & Glauner (2022) broaden these equivalence findings to encompass all risk-aware MDPs that employ spectral risk measures, which are known as a class of coherent risk measures like the Expected Shortfall. However, the risk measures examined are confined to the Markov risk measure (Ruszczyński 2010), known to be potentially non-gradient-dominant (Huang et al., 2021).
@@
where $Q_{\tau}(X):=\inf\{x\mid P(X\leq x)\geq\tau\}$ and $\tau\in(0,1)$ is the quantile value. The authors reformulate (60) as a max-min optimization and have proven that the quantile-based value-to-go function is equivalent to solving a specific optimization problem for each action. This breakthrough facilitates the application of dynamic programming techniques. Additionally, they craft efficient algorithms and conduct a detailed complexity analysis. In terms of model-free approaches, Yu & Shen (2022) adapt CVaR as a special one-step conditional risk measure to preserve time consistency and subsequently reformulate the risk-averse MDP as a risk-neutral counterpart with augmented action space and the adjusted immediate rewards. Furthermore, the authors show that the Bellman operator in this formulation is a contraction mapping, and consequently, the authors extend Q-learning results into this case and develop a risk-averse deep Q-learning algorithm.
The concept of weaving regularization into dynamic programming also has been thoroughly investigated within the reinforcement learning field. (Kaufman & Schaefer 2013, Neu et al. 2017, Kostrikov et al. 2021, Ho et al. 2021, Kumar et al. 2022). While these regularized variants of value iteration and policy iteration have been developed, Grand-Clément & Petrik (2022) is the first to propose the convex optimization formulation of RMDPs. This innovative approach integrates an entropic regularization of the robust Bellman operator with a change of variables involving exponential and logarithmic functions. By conducting this convex reformulation, arbitrary algorithms designed for solving convex optimization problems can be incorporated with the reformulated RMDPs. This novel perspective has paved the way for the development of innovative algorithms tailored to RMDPs.
@@ 8 Conclusion
This survey provides an overview of the theory, methodologies, and advances of RMDPs with uncertain transition probability. Before the details of the literature on RMDPs, we elaborate on the foundations of RMDPs and DRMDPs and introduce an important concept rectangularity. Rectangularity plays a key role in most RMDPs research, allowing decomposability to ensure that RMDPs can be solved in polynomial time. We summarize three class definitions of rectangularity and propose a new and straightforward proof that solving non-rectangular RMDPs is NP-hard.
Mainly, we categorize the ambiguity modeling approaches into three groups, parametric, moment-based, and discrepancy-based. The parametric approach is the simplest yet most limited way among the three methods to characterize the uncertainty. Since it usually imposes individual support constraints or limits the distribution family, it exhibits statistically poor performance and heavily relies on LPs to compute the optimal policy. Due to the issues of interpretability, moment-based ambiguity modeling is more suitable for situations where the randomness is exogenous or concrete problem backgrounds. As the increasingly popular approach, discrepancy-based modeling enjoys desirable statistical guarantees and performance in a data-driven manner. We summarize three prevalent discrepancy measures–norm, $\phi$-divergence, and Wasserstein distance. All of them provide powerful finite sample guarantees to contain the underlying probability measure within constructed ambiguity sets with (high) probability, in turn, guiding how many samples we need to make a safe decision. It should not be ignored that the research of fast algorithms based on discrepancy-based ambiguity sets has gained more and more attention.
Finally, we review recent efforts that depart from rectangular assumptions and minimax framework. Relaxing traditional rectangular assumptions allows capturing of dependencies between transitions, which better characterizes the reasonable impact of uncertainty and reduces conservatism. However, these coupled uncertainty models often need to compromise flexibility and generality. Besides the minimax objective, there are also other approaches to hedge uncertainty and achieve robust performance. These new paradigms bring new techniques and perspectives, complementing the research of RMDPs.
Sequential decision-making under uncertainty is a difficult yet crucial research problem in practice. Although RMDPs have a substantial theoretical basis, their applications in sequential decision-making are relatively limited compared to other tools. One potential explanation for this limitation is rectangularity. How to reasonably relax this assumption and develop new frameworks is an important direction, particularly within a data-driven context and when integrating artificial intelligence techniques. Meanwhile, the development of fast algorithms and the provably robust RL algorithms will be an interesting and practical direction.